# CORE: Mitigating Catastrophic Forgetting in Continual Learning through Cognitive Replay

### 02 Feb 2024 v1

@@ Abstract
This paper introduces a novel perspective to significantly mitigate catastrophic forgetting in continuous learning (CL), which emphasizes models’ capacity to preserve existing knowledge and assimilate new information. Current replay-based methods treat every task and data sample equally and thus can not fully exploit the potential of the replay buffer. In response, we propose COgnitive REplay (CORE), which draws inspiration from human cognitive review processes. CORE includes two key strategies: Adaptive Quantity Allocation and Quality-Focused Data Selection. The former adaptively modulates the replay buffer allocation for each task based on its forgetting rate, while the latter guarantees the inclusion of representative data that best encapsulates the characteristics of each task within the buffer. Our approach achieves an average accuracy of 37.95% on split-CIFAR10, surpassing the best baseline method by 6.52%. Additionally, it significantly enhances the accuracy of the poorest-performing task by 6.30% compared to the top baseline.
@@ Introduction
In response to the catastrophic forgetting in continual learning, current methodologies predominantly fall into three categories: Parameter Isolation, Regularization, and Data Replay (?, ?).
@@
Forgetting in models parallels human cognitive processes, notably cognitive overload and interference. “Cognitive overload” in models occurs when the fixed size of the model, including its architecture and parameter capacity, becomes saturated with new information, hindering its ability to effectively retain and process previously learned information (?, ?). This resembles human cognitive constraints under substantial information loads. “Interference”, inherent in the neural network’s interconnected architecture, means even minor parameter adjustments can drastically impact task recall (?, ?). This phenomenon reflects interference-based forgetting in human memory, where new learning can obstruct the retention of established knowledge.
@@
Comprehensive Performance Evaluation Across Baselines
@@
Analysis of Per-Task Precision in Progressive Learning Scenarios

### 09 Apr 2024 v2

@@ Abstract
This paper introduces a novel perspective to significantly mitigate catastrophic forgetting in continuous learning (CL), which emphasizes models’ capacity to preserve existing knowledge and assimilate new information. Current replay-based methods treat every task and data sample equally and thus can not fully exploit the potential of the replay buffer. In response, we propose COgnitive REplay (CORE), which draws inspiration from human cognitive review processes. CORE includes two key strategies: Adaptive Quantity Allocation and Quality-Focused Data Selection. The former adaptively modulates the replay buffer allocation for each task based on its forgetting rate, while the latter guarantees the inclusion of representative data that best encapsulates the characteristics of each task within the buffer. Our approach achieves an average accuracy of 37.95% on split-CIFAR10, surpassing the best baseline method by 6.52%. Additionally, it significantly enhances the accuracy of the poorest-performing task by 6.30% compared to the top baseline. Code is available at https://github.com/sterzhang/CORE.
@@ Introduction
In response to the catastrophic forgetting in continual learning, current methods predominantly fall into three categories: Parameter Isolation, Regularization, and Data Replay.
@@
Forgetting in models parallels human cognitive processes, notably cognitive overload and interference. Cognitive overload in models occurs when the fixed size of the model, including its architecture and parameter capacity, becomes saturated with new information, hindering its ability to effectively retain and process previously learned information. This resembles human cognitive constraints under substantial information loads. Interference, inherent in the neural network’s interconnected architecture, means even minor parameter adjustments can drastically impact task recall (?, ?). This phenomenon reflects interference-based forgetting in human memory, where new learning can obstruct the retention of established knowledge.
@@
Evaluation Across Baselines
@@
Analysis of Progressive Learning Scenarios
@@ Conclusion
Acknowledgments
This research was supported in part by the National Natural Science Foundation of China under grant No. 62172303.